What is the total surface area of a rectangular prism box that measures 5 feet by 1 foot by 1 foot?

Answer by checkley77(12569)

2(1*1)+4(5*1)=2*1=4*5=2+20=22 ft^2 is the totala surface area of the box.
[ends] [top,bottom & 2 sides]